OpenStudy (anonymous):

Verify the identity. cot(x-pi/2)=-tanx

OpenStudy (zzr0ck3r):

didnt we already do this?

OpenStudy (anonymous):

I believe so, yet I'm still ??????????? about this new one that came at me. Sorry if I annoyed you...

OpenStudy (zzr0ck3r):

\(\cos(x-\frac{\pi}{2})=\sin(x)\\\sin(x-\frac{\pi}{2})=-\cos(x)\) So \(\cot(x-\frac{\pi}{2}) = \frac{\cos(x-\frac{\pi}{2})}{sin(x-\frac{\pi}{2})}=\frac{\sin(x)}{-\cos(x)}=-\frac{\sin(x)}{\cos(x)}=-\tan(x)\)

OpenStudy (zzr0ck3r):

just thought it was crazy if two people asked the same exact question

OpenStudy (zzr0ck3r):

if you want to verify the first two lines, use \(\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)\) and \(\sin(A-B)=\sin(A)\cos(B)-\sin(B)\cos(A)\)

OpenStudy (anonymous):

//I hate this class so much omg// :) Thanks!

OpenStudy (zzr0ck3r):

example \(\cos(x-\frac{\pi}{2})=\cos(x)\cos(\frac{\pi}{2})+\sin(x)\sin(\frac{\pi}{2})=\cos(x)*0+\sin(x)*1=\sin(x)\)

OpenStudy (zzr0ck3r):

np